# Gradient flow in parameter space is equivalent to linear interpolation in output space

### v2

@@ 1. Introduction
We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank deficient matrices is provided as well.
@@ 2. Main results
$\left\{\underline{\theta}:\nabla_{\underline{x}(\underline{\theta})}\mathcal{C}=0\right\}\subset\left\{\underline{\theta}:\nabla_{\underline{\theta}}\mathcal{C}=0\right\},$ (2.16)
however, the two sets are not necessarily identical. It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$. Moreover, if $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss. See e.g. [Coo18, TKB19, Che23, KMTM24, CE25] for more details.
Notation. Given a matrix $A\in\mathbb{R}^{m\times n}$, we denote by $(A){}^{+}$ its pseudoinverse, or Moore–Penrose inverse; see appendix A for a definition and useful properties. Moreover, given a subspace $V\subset\mathbb{R}^{d}$, we denote by $\mathcal{P}_{V}$ the orthogonal projection onto $V$, such that $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$.
@@ 2.1. Adapted gradient flow
$DD^{T}\psi=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.21)
@@
$\partial_{s}\underline{x}(s)=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.22)
@@
$=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$
@@
$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{ran}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
We will refer to (2.26) as adapted gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ and (2.27) as constrained Euclidean gradient flow in $\mathbb{R}_{\underline{x}}^{QN}$.
@@
This holds for any loss function $\ell$ in (2.7), and to any neural network whose activation function has Lipschitz continuous derivatives. Non-smooth activation functions, such as ReLU, can be mollified to satisfy this condition.
@@
For $\alpha\in[0,1)$, note that $A_{\underline{\theta},\alpha}>0$ for all $\underline{\theta}$. For $\alpha=1$, writing $\mathcal{P}_{\operatorname{ran}(D^{T})}=(D){}^{+}D$, we have
@@
since (2.15) shows $\nabla_{\underline{\theta}}\mathcal{C}\in\operatorname{ran}(D^{T})$, and also
$\ker(D^{T}D){}^{+}=\ker(D^{T}D)^{T}=\ker(D^{T}D)=\ker((D){}^{+}D),$ (2.35)
@@
=$\stackrel{{\scriptstyle(*)}}{{=}}-\left(\alpha\mathcal{P}_{\operatorname{ran}(DD^{T})}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$
$=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$
where ()(*)( ∗ ) follows from (A.7), (A.8) and (A.12). Therefore,
@@
$V_{\underline{x}(s),\alpha}:=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)D(s)D(s)^{T}\right)\mathcal{P}_{\operatorname{ran}(D(s)D(s)^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.40)
@@
$V_{\underline{x}(s),1}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.42)
so that in fact $V_{\underline{x}(s),\alpha},\,\alpha\in[0,1],$ interpolates between the standard dynamics given by (2) at $\alpha=0$, and constrained Euclidean gradient flow (2.27) for $\underline{x}(s)$ at $\alpha=1$. Moreover, for $\alpha\in(0,1]$, $(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T})$ is positive-definite, and for $\alpha=0$, $DD^{T}$ is injective in the range of $\mathcal{P}_{\operatorname{ran}(DD^{T})}$, so
$(\alpha{\textbf{1}}+(1-\alpha)DD^{T})\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0\text{ if, and only if, }\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0.$ (2.43)
@@
$[h_{\alpha}]=\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\Big{|}_{\operatorname{ran}(DD^{T}[\underline{\theta}])}\,.$ (2.45)
@@ 2.2. Reparametrization for the $L^{2}$ loss
where $\underline{\tilde{x}}(t)=\underline{x}[\underline{\theta}(s(t))]$ and $\mathcal{P}_{t}:=\mathcal{P}_{\operatorname{ran}(DD^{T})}[\underline{\theta}(s(t))]$. When $\operatorname{rank}(D)<QN$, the deviation from linear interpolation is given by
[…]
@@
Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula, and substituting the resulting expression in (2.60) yields (2.58), as claimed. ∎
@@ 2.3. Cross-entropy loss
where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$.
@@
[…]
@@
and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$111 While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work. This scenario can occur when certain training inputs do not clearly belong to a single class, or can be forced by subjecting the labels to small perturbations. We have thus proved most of the following:
@@ 3.1. Prescribed paths in output space
$\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}=(D){}^{+}\partial_{s}\underline{x},$ (3.1)
and so assuming $\partial_{s}\underline{\theta}=\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}$, [^2] Note that this is satisfied by the standard gradient flow $\partial_{s}\underline{\theta}=-\nabla_{\underline{\theta}}\mathcal{C}=-D^{T}\nabla_{\underline{x}}\mathcal{C}.$ we can write
@@
$\partial_{s}\underline{x}(s)$ $\in\operatorname{ran}(D(s)),$ (3.3)
@@
$\operatorname{ran}(D)=(\ker(D^{T}))^{\perp}=(\ker(D){}^{+})^{\perp},$ (3.6)
@@
$(\underline{x}_{0}-\underline{y})\in\operatorname{ran}{D(s)},$ (3.7)
@@
Whether or not a path $\underline{x}(s)$ satisfies the constraint in (3.1) (or (3.7) in the example above) could depend on the choice of $\underline{\theta}_{0}$. In general, properties of the Jacobian matrix $D[\underline{\theta}(s)]$ that determine the dynamics of $\underline{x}(s)$, such as rank and the subspace $\operatorname{ran}(DD^{T})$, might depend on the initialization even within a class $\left\{\underline{\theta}_{0}:\underline{x}(\underline{\theta}_{0})=\underline{x}_{0}\right\}$.
@@ 3.3. Tangent kernel
From (3.15) we see that his holds if, and only if, $DD^{T}$ is a positive-definite matrix.
@@ 4. Some related work
Certain works study the rank of $D$ more explicitly: In [GLMW22], the authors study the functional dimension of ReLU neural networks, defined in terms of the rank of the Jacobian. In [KMTM24], the authors show that for mildly overparametrized ReLU neural networks with scalar output ($Q=1$) most activation patterns correspond to regions in parameter space where the Jacobian is full rank.
Acknowledgments: T.C. gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE. P.M.E. was supported by NSF grant DMS-2009800 through T.C., and UT Austin’s GS Summer fellowship. T.C. and P.M.E thank the anonymous referee who suggested we investigate the cross-entropy loss.
@@ Appendix A Pseudoinverse
$A(A){}^{+}$ $=\mathcal{P}_{\operatorname{ran}(A)},$ (A.8)
$(A){}^{+}A$ $=\mathcal{P}_{\operatorname{ran}(A^{T})}.$ (A.9)
@@
$\operatorname{ran}(A){}^{+}$ $=\operatorname{ran}A^{T},$ (A.10)

### v3

@@ 1. Introduction
These results underscore the central role of the rank of the Jacobian (or equivalently, the neural tangent kernel) in shaping optimization dynamics. In particular, if the gradient flow for a convex loss in output space reaches a fixed point, then at least one of the following must be true: (i) the solution interpolates the data, or (ii) the Jacobian is not injective. This becomes evident when considering the constrained Euclidean gradient flow in output space, as is done for the $L^{2}$ loss in section 2.2 and for the cross-entropy loss in section 2.3.
@@
We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank-deficient matrices is provided as well.
@@ 2. Main results
$\nabla_{\underline{x}(\underline{\theta})}\mathcal{C}=0\implies\nabla_{\underline{\theta}}\mathcal{C}=0,$ (2.16)
but the converse does not hold in general. It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0\iff\nabla_{\underline{\theta}}\mathcal{C}=0$. If $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss. See e.g. [TKB19, Coo21, Che23, KMTM24, CE25].
More generally, consider a cost function for which the only minima in output space are global minima where the model has interpolated the data, $f(\underline{\theta}_{*},x)=y(x)$ for all $x\in\mathcal{X}_{0}$; this is the case for the $L^{2}$ and cross-entropy losses. Then for an overparametrized model ($K\geq QN$), we can read the following from (2.15): If the corresponding gradient flow reaches a fixed point, then at least one of the following holds,
  • the model interpolates the data, or
  • $D$ is not full-rank.
In other words, if the interpolating solution is not attainable, gradient flow will only converge when $D$ loses rank.
Notation. Given a matrix $A\in\mathbb{R}^{m\times n}$, we denote by $(A){}^{+}$ its pseudoinverse, or Moore–Penrose inverse; see appendix A for a definition and useful properties. Moreover, given a subspace $V\subset\mathbb{R}^{d}$, we denote by $\mathcal{P}_{V}$ the orthogonal projection onto $V$, such that $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$. In this work, $V$ will usually be the range (or image) of a linear operator.
@@ 2.1. Adapted gradient flow
$DD^{T}\psi=\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.21)
@@
$\partial_{s}\underline{x}(s)=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.22)
@@
$=\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$
@@
$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{range}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
We will refer to (2.26) as adapted gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ and (2.27) as constrained Euclidean gradient flow in $\mathbb{R}_{\underline{x}}^{QN}$. [^1] Note that this is not the same as the Projected Gradient Descent algorithm for optimization on a feasible set. In (2.27) the direction of the flow is constrained, not the value of $\underline{x}$.
@@
This holds for any loss function $\ell$ in (2.7), and to any neural network whose activation function has Lipschitz continuous derivatives. Commonly used non-smooth activation functions, such as ReLU, can be smoothed to satisfy this condition, although the resulting gradient flow might depend on the smoothing.
@@
For $\alpha\in[0,1)$, note that $A_{\underline{\theta},\alpha}>0$ for all $\underline{\theta}$. For $\alpha=1$, writing $\mathcal{P}_{\operatorname{range}(D^{T})}=(D){}^{+}D$, we have
@@
since (2.15) shows $\nabla_{\underline{\theta}}\mathcal{C}\in\operatorname{range}(D^{T})$, and also
[…]
@@
$\stackrel{{\scriptstyle(*)}}{{=}}-\left(\alpha\mathcal{P}_{\operatorname{range}(DD^{T})}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$
$=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$
where ()(*) follows from (A.7), (A.8) and (A.12). Therefore,
@@
$V_{\underline{x}(s),\alpha}:=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)D(s)D(s)^{T}\right)\mathcal{P}_{\operatorname{range}(D(s)D(s)^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.40)
@@
$V_{\underline{x}(s),1}=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.42)
so that in fact $V_{\underline{x}(s),\alpha},\,\alpha\in[0,1],$ interpolates between the standard dynamics given by (2) at $\alpha=0$, and constrained Euclidean gradient flow (2.27) for $\underline{x}(s)$ at $\alpha=1$. Moreover, for $\alpha\in(0,1]$, $(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T})$ is positive-definite, and for $\alpha=0$, $DD^{T}$ is injective in the range of $\mathcal{P}_{\operatorname{range}(DD^{T})}$, so
$(\alpha{\textbf{1}}+(1-\alpha)DD^{T})\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0\text{ if, and only if, }\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0.$ (2.43)
@@
$[h_{\alpha}]=\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\Big|_{\operatorname{range}(DD^{T}[\underline{\theta}])}\,.$ (2.45)
@@ 2.2. Reparametrization for the $L^{2}$ loss
Remark 2.4.
This implies that $\underline{y}$ is in the image of $\underline{x}$, which is not necessarily the case in general. Recall from Remark 2.1 that if the interpolating solution is not attainable, then $D$ must lose rank to reach a fixed point. Equation (2.49) only holds if $D(t)$ is full rank for all $t\in[0,s]$ and the initialization $\underline{\theta}(0)$.
@@
where $\underline{\tilde{x}}(t)=\underline{x}[\underline{\theta}(s(t))]$ and $\mathcal{P}_{t}:=\mathcal{P}_{\operatorname{range}(DD^{T})}[\underline{\theta}(s(t))]$. When $\operatorname{rank}(D)<QN$, the deviation from linear interpolation is given by
$\underline{\tilde{x}}(t)-\big((1-t)\underline{x}_{0}+t\underline{y}\big)=\int_{0}^{t}dt^{\prime}\,\mathcal{U}_{t,t^{\prime}}\,\frac{1-t}{1-t^{\prime}}\,\mathcal{P}^{\perp}_{t^{\prime}}\,(\underline{x}_{0}-\underline{y}),$ (2.58)
@@
Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula (see e.g., [Tes12]), and substituting the resulting expression in (2.60) yields (2.58), as claimed. ∎
@@ 2.3. Cross-entropy loss
Assuming $\operatorname{rank}{D}=QN\leq K$, the gradient flow in output space resulting from (2.26) is simply Euclidean gradient flow,
@@
where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$. This is in line with the fact that the softmax function is invariant under addition of a scalar multiple of $u_{Q}$.
@@
$D^{2}\ell\,\rvert_{f_{t}(x)}\,u_{Q}=\sigma(f_{t}(x))-\sigma(f_{t}(x))\underbrace{(\sigma(f_{t}(x))\cdot u_{Q})}_{=\sum_{j}\sigma(f_{t}(x))_{j}\,=\,1}=0.$ (2.77)
@@
and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$. [^2] While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work. This scenario can occur when certain training inputs do not clearly belong to a single class, or can be forced by subjecting the labels to small perturbations. We have thus proved most of the following:
@@
Remark 2.9.
As in the previous section, this corresponds to an interpolating solution. See Remarks 2.1 and 2.4.
@@ 3.1. Prescribed paths in output space
$\mathcal{P}_{\operatorname{range}(D^{T})}\partial_{s}\underline{\theta}=(D){}^{+}\partial_{s}\underline{x},$ (3.1)
and so assuming $\partial_{s}\underline{\theta}=\mathcal{P}_{\operatorname{range}(D^{T})}\partial_{s}\underline{\theta}$, [^3] Note that this is satisfied by the standard gradient flow $\partial_{s}\underline{\theta}=-\nabla_{\underline{\theta}}\mathcal{C}=-D^{T}\nabla_{\underline{x}}\mathcal{C}.$ we can write
@@
$\partial_{s}\underline{x}(s)$ $\in\operatorname{range}(D(s)),$ (3.3)
@@
$\operatorname{range}(D)=(\ker(D^{T}))^{\perp}=(\ker(D){}^{+})^{\perp}$, (3.6)
@@
$(\underline{x}_{0}-\underline{y})\in\operatorname{range}{D(s)},$ (3.7)
@@
Whether or not a path $\underline{x}(s)$ satisfies the constraint in (3.1) (or (3.7) in the example above) could depend on the choice of $\underline{\theta}_{0}$. In general, properties of the Jacobian matrix $D[\underline{\theta}(s)]$ that determine the dynamics of $\underline{x}(s)$, such as rank and the subspace $\operatorname{range}(DD^{T})$, might depend on the initialization even within a class $\left\{\underline{\theta}_{0}:\underline{x}(\underline{\theta}_{0})=\underline{x}_{0}\right\}$.
@@ 3.3. Tangent kernel
From (3.15) we see that this holds if, and only if, $DD^{T}$ is a positive-definite matrix.
@@ 4. Some related work
Preconditioners
The preconditioner $(D^{T}D)^{+}$ introduced here closely resembles the inverse of the generalized Gauss–Newton matrix (GN), which approximates the Hessian of the loss with respect to the parameters:
$\left(D^{T}D\right)_{ij}$ $=\,\,\,\,(\partial_{i}\underline{x}^{T})(\partial_{j}\underline{x}),$
$\frac{\partial^{2}\mathcal{C}}{\partial\theta_{i}\partial\theta_{j}}\,\,\,$ $=\underbrace{(\partial_{i}\underline{x}^{T})H_{\underline{x}}(\partial_{j}\underline{x})}_{(GN)_{ij}}+\partial_{i}\partial_{j}\underline{x}^{T}\nabla_{\underline{x}}\mathcal{C}.$ (4.1)
The distinction lies in omitting $H_{\underline{x}}$, the Hessian with respect to the outputs, which is the identity for the $L^{2}$ loss. The proof of Theorem 2.3 remains valid for $GN^{-1}$ as long as $H_{\underline{x}}$ is positive-semidefinite in the region of interest (which holds for the $L^{2}$ and cross-entropy losses), but fails in general. Recent work has examined the properties of the GN matrix [ZSL24], as well as potential advantages [AVKM25] of the full Gauss–Newton matrix compared to Adam [KB17]. For a broader discussion on curvature-aware optimization, see e.g. [Mar20] and references therein.
@@
Certain works study the rank of $D$ more explicitly: In [GLMW25], the authors study the functional dimension of ReLU neural networks, defined in terms of the rank of the Jacobian. In [KMTM24], the authors show that for mildly overparametrized ReLU neural networks with scalar output ($Q=1$) most activation patterns correspond to regions in parameter space where the Jacobian is full rank.
Acknowledgments: T.C. gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE. P.M.E. was supported by NSF grant DMS-2009800 through T.C., and UT Austin’s GS Summer fellowship. T.C. and P.M.E. thank the anonymous referee who suggested we investigate the cross-entropy loss.
@@ Appendix A Pseudoinverse
$A(A){}^{+}$ $=\mathcal{P}_{\operatorname{range}(A)},$ (A.8)
$(A){}^{+}A$ $=\mathcal{P}_{\operatorname{range}(A^{T})}.$ (A.9)
@@
$\operatorname{range}(A){}^{+}$ $=\operatorname{range}A^{T},$ (A.10)